OpenStudy (nilankshi):

A capsule of medicine is in the shape of a sphere of diameter 3.5 mm.how much medicine (in nm^3 )is needed to fill this capsule?

OpenStudy (nilankshi):

help mee plzz.....

OpenStudy (phi):

can you find the volume in cubic mm?

OpenStudy (nilankshi):

yes.

OpenStudy (phi):

1 cubic nm = 1 nm^3 = 1e^-27 m^3 1 mm= 1e-3 m and 1 mm^3 (cubic mm) = 1e-9 m^3

OpenStudy (sasogeek):

volume of a sphere is \[\huge \frac{4}{3}\pi r^3 \] diameter is 3.5, implying that radius is 1.75, substitute and u have a value in mm^3, then convert to nm^3

OpenStudy (anonymous):

22.458 mm^3

OpenStudy (anonymous):

What factors cause mechanical weathering? Explain why chemical weathering occurs faster in hot, wet climates than in cool, dry climates. What role does weathering play in the formation of soil? How did settlers on the Great Plains help create the Dust Bowl? Describe two techniques that farmers use to conserve soil.

OpenStudy (phi):

so I would do: \[\frac{4}{3} \pi (3.5)^3 \cdot 10^{-9} m^3\cdot \frac{1 nm^3}{10^{-27} m^3}\]

OpenStudy (sasogeek):

phi, the diameter is 3.5, not the radius... :|

OpenStudy (phi):

* 3.5/2

OpenStudy (phi):

tweaked \[ \frac{4}{3} \pi (3.5/2)^3 \cdot 10^{-9} m^3\cdot \frac{1 nm^3}{10^{-27} m^3} \]